# Procedure Learning via Regularized Gromov-Wasserstein Optimal Transport

### v1

@@ Abstract
We study the problem of self-supervised procedure learning, which discovers key steps and establishes their order from a set of unlabeled procedural videos. Previous procedure learning methods typically learn frame-to-frame correspondences between videos before determining key steps and their order. However, their performance often suffers from order variations, background/redundant frames, and repeated actions. To overcome these challenges, we propose a self-supervised procedure learning framework, which utilizes a fused Gromov-Wasserstein optimal transport formulation with a structural prior for computing frame-to-frame mapping between videos. However, optimizing exclusively for the above temporal alignment term may lead to degenerate solutions, where all frames are mapped to a small cluster in the embedding space and hence every video is associated with only one key step. To address that limitation, we further integrate a contrastive regularization term, which maps different frames to different points in the embedding space, avoiding the collapse to trivial solutions. Finally, we conduct extensive experiments on large-scale egocentric (i.e., EgoProceL) and third-person (i.e., ProceL and CrossTask) benchmarks to demonstrate superior performance by our approach against previous methods, including OPEL which relies on a traditional Kantorovich optimal transport formulation with an optimality prior.
@@ 1 Introduction
Understanding how to perform complex tasks by watching others is a hallmark of human intelligence. An increasingly important goal in computer vision is to replicate this ability in machines, enabling them to learn a procedure from watching video demonstrations. Procedure learning (PL) refers to the task of identifying key steps and their ordering from instructional videos [2, 41, 74]. Unlike action recognition and segmentation, which focus on analyzing single videos and classifying short-term activities [62, 10, 55], procedure learning extracts semantics such as key steps and their order across videos, even when step appearances vary, their order differ, or background/redundant frames occur frequently, as shown in Fig. 1. These challenges are especially common in real-world datasets, where videos exhibit non-monotonic step sequences, repeated actions, and untrimmed content that results in irrelevant frames.
@@
While earlier methods relied on full or weak supervision [7, 33] which require substantial human effort and do not scale to large task domains, recent works have focused on self-supervised alignment of instructional videos. CnC [4], for instance, uses temporal cycle-consistency and contrastive learning to align steps across videos. However, it is sensitive to step order variation and background/redundant frames. More recent works have employed optimal transport (OT) formulations for alignment due to their flexibility. One such method, VAVA [47], frames alignment as an OT problem to accommodate variations in step ordering and incorporates an optimality prior along with inter-video and intra-video contrastive losses to solve the temporal misalignment problem. However, a key limitation in VAVA [47] is that it is challenging to balance multiple losses and handle repeated actions, as noted by [17, 3]. Recently, OPEL [11] applies VAVA for self-supervised procedure learning, inheriting these limitations.
@@
In this paper, we propose a self-supervised procedure learning framework built upon a fused Gromov-Wasserstein optimal transport (FGWOT) formulation with a structural prior. Unlike previous OT-based approaches [47, 11] that utilize a traditional Kantorovich OT formulation, our method not only aligns frames based on appearance but also enforces a structural prior (i.e., temporal consistency) across videos. This makes our model better suited for handling step order variations, background/redundant frames, and repeated actions. Moreover, we empirically identify that optimizing only for temporal alignment may lead to degenerate solutions, where all frames collapse to a single cluster in the embedding space. To prevent this issue, we utilize Contrastive Inverse Difference Moment (C-IDM) as a regularization that encourages embedding diversity across frames. Our regularized Gromov-Wasserstein optimal transport (RGWOT) approach uses a unified loss with a purpose-aligned regularization, avoiding the difficulty of balancing multiple losses and conflicting regularizations. Our model outperforms prior works on both egocentric [4] and third-person [23, 74] procedure learning benchmarks..
@@
  • We introduce an optimal transport-based framework for self-supervised procedure learning, which adopts a fused Gromov-Wasserstein optimal transport formulation with a structural prior for establishing frame-to-frame correspondences between videos.
@@
  • Extensive evaluations on egocentric (i.e., EgoProceL) and third-person (i.e., ProCeL and CrossTask) datasets show that our approach outperforms previous methods, including OPEL which adopts a classical Kantorovich optimal transport formulation with an optimality prior.
@@ 2 Related Work
Self-Supervised Learning. Self-supervised learning uses pretext tasks that generate a supervision signal from the data itself. Early self-supervised approaches focused mainly on images, leveraging spatial cues through tasks such as image colorization [43, 44], object counting [48, 52], solving puzzles [9, 38, 37], predicting rotations [24, 27], and image inpainting [36]. Recently, self-supervised learning has expanded to video data, which inherently offers spatial and temporal structures. Video-based pretext tasks include forecasting future frames [1, 16, 63, 67, 30], enforcing temporal coherence [28, 51, 75], predicting the temporal order of frames [25, 45, 50, 70], determining the arrow of time [54, 69], predicting the pace of actions [6, 68, 72], and frame clustering [42, 65]. These tasks help models capture motion and temporal cues, yielding richer representations for downstream tasks. In this work, we focus on self-supervised procedure learning for identifying key steps and their order from videos.
@@
Procedure Learning. Many procedure learning (PL) methods focus on learning frame-level features that capture task structure [23, 41, 22, 66, 4]. For example, Kukleva et al. [41] improve the representation by using relative frame timestamps, while VidalMata et al. [66] focus on predicting future frames and their timestamps. Elhamifar et al. [22] utilize attention mechanism on individual frames for better feature learning. Bansal et al. [4] use temporal correspondences across videos to create signals for robust frame-level embeddings. Recently, Bansal et al. [5] use a task-level graph representation to cluster frames that are semantically similar and temporally close. Beyond purely visual approaches, procedure learning has also been explored in multi-modal settings, combining narrated text and videos [14, 73, 49, 2, 18, 26], as well as modalities like optical flow, depth, and gaze information [58]. These multi-modal approaches generally assume reliable alignment between video content and the accompanying modalities [2, 49, 73], an assumption that frequently fails due to asynchrony among data streams [23, 22]. Also, reliance on imperfect automatic speech recognition systems requires manual correction, and multi-modal integration adds memory and computational overhead. Recent model for egocentric PL [11] focuses exclusively on visual data, bypassing the inaccuracies associated with multi-modal alignment. Here, we propose an advanced optimal transport formulation, outperforming [11] on visual procedure learning.
@@
Video Alignment. TCC [21] enforces cycle consistency between frames, while GTCC [17] extends this with multi-cycle consistency to better handle repeated actions. However, both methods focus on local alignment and do not capture the global temporal structure. For global alignment, LAV [31] applies dynamic time warping [13] but assumes monotonic ordering without background/redundant frames. VAVA [47] addresses non-monotonic sequences and background/redundant frames using a classical Kantorovich optimal transport formulation [12] but suffers from balancing multiple losses and repeated actions. Recently, VAOT [3] develops a fused Gromov-Wasserstein optimal transport formulation which can handle order variations, background/redundant frames, and repeated actions. In this paper, we empirically show a possible degenerate solution for VAOT [3] and propose a regularized optimal transport formulation to address that issue.
@@
Learning Key-Step Ordering. People often perform the same task using different orders of key-steps, but most previous methods fail to model this variation. Many approaches either assume a strict step ordering [23, 41, 66] or do not predict the order at all [22, 59]. To better capture how tasks are actually completed, it is important to account for these different key-step sequences. Our method addresses this by creating a specific sequence of key-steps for each video, allowing the model to infer the particular order required to complete the task.
@@ 3 Our Approach
We present our self-supervised procedure learning pipeline, which includes finding frame-to-frame correspondences between videos via a regularized Gromov-Wasserstein optimal transport formulation in Sec. 3.1 and performing clustering and key-step ordering in Sec. 3.2.
Notations. Let us represent two input videos of $N$ and $M$ frames as $X=\{x_{1},\dots,x_{N}\}$ and $Y=\{y_{1},\dots,y_{M}\}$ respectively. The frame-level embeddings of $X$ and $Y$ are then written as $\mathbf{X}=f_{\bm{\theta}}(X)\in\mathbb{R}^{N\times D}$ and $\mathbf{Y}=f_{\bm{\theta}}(Y)\in\mathbb{R}^{M\times D}$ respectively, where $f_{\bm{\theta}}$ (with learnable parameters $\bm{\theta}$) is the embedding function and $D$ is the length of the embedding vector. $K$ is the number of key steps. In addition, the dot product of $\mathbf{A},\mathbf{B}\in\mathbb{R}^{n\times m}$ is expressed as $\langle\mathbf{A},\mathbf{B}\rangle=\sum_{i,j}A_{ij}B_{ij}$, while a vector of ones is written as $\mathbf{1}_{n}\in\mathbb{R}^{n}$. Finally, $[n]=\{1,\dots,n\}$ models a discrete set of $n$ elements. We represent the $(m-1)$ dimensional probability simplex as $\Delta_{m}\subset\mathbb{R}^{m}$ and the Cartesian product space consisting of $n$ such simplexes as $\Delta_{m}^{n}\subset\mathbb{R}^{m\times n}$.
@@ 3.1.1 Fused Gromov-Wasserstein Optimal Transport
We consider temporal video alignment as an optimal transport problem in the discrete setting. We follow [3] to apply a fused Gromov-Wasserstein optimal transport formulation (FGWOT), which is a combination of Kantorovich optimal transport (KOT) and Gromov-Wasserstein optimal transport (GWOT). Below we provide the details of KOT and GWOT. In particular, given a ground cost $\mathbf{C}\in\mathbb{R}^{n\times m}_{+}$, the KOT objective seeks the minimum-cost coupling $\mathbf{T}^{\star}$ between histograms $\mathbf{p}\in\Delta_{n}$ and $\mathbf{q}\in\Delta_{m}$ as:
@@
Cost Matrices. The cost matrix $\mathbf{C}\in\mathbb{R}^{N\times M}_{+}$ in the KOT component measures visual differences between $X$ and $Y$ and is defined as $\mathbf{C}_{ij}=1-\frac{\mathbf{x}_{i}^{\top}\mathbf{y}_{j}}{\|\mathbf{x}_{i}\|_{2}\|\mathbf{y}_{j}\|_{2}}$, with frame embeddings $\mathbf{x}_{i}=f_{\bm{\theta}}(x_{i})$ and $\mathbf{y}_{j}=f_{\bm{\theta}}(y_{j})$. Also, we follow [47] to augment $\mathbf{C}$ as $\tilde{\mathbf{C}}=\mathbf{C}+\rho\mathbf{R}$, where $\rho\geq 0$ is a balancing parameter and the temporal prior $\mathbf{R}$ (with $\mathbf{R}_{ij}=\left|i/N-j/M\right|$) enforces the coupling $\mathbf{T}$ to have a banded diagonal shape. Moreover, we set $L(a,b)=ab$, and follow [3] to define cost matrices $\mathbf{C}^{x}\in\mathbb{R}_{+}^{N\times N}$ and $\mathbf{C}^{y}\in\mathbb{R}_{+}^{M\times M}$ in the GWOT component as:
@@
with $\delta_{ik}=|i-k|$, $\delta_{jl}=|j-l|$, and a radius parameter $r\in(0,1]$. The GWOT component imposes temporal consistency over $\mathbf{T}$. Specifically, mapping temporally adjacent frames in $X$ ($\delta_{ik}\leq Nr$) to temporally remote frames in $Y$ ($\delta_{jl}>Mr$) induces a cost ($L(\mathbf{C}^{x}_{ik},\mathbf{C}^{y}_{jl})=\frac{1}{r}$), however assigning temporally neighboring frames in $X$ ($\delta_{ik}\leq Nr$) to temporally nearby frames in $Y$ ($\delta_{jl}\leq Mr$) or assigning temporally remote frames in $X$ ($\delta_{ik}>Nr$) to temporally distant frames in $Y$ ($\delta_{jl}>Mr$) induces no cost ($L(\mathbf{C}^{x}_{ik},\mathbf{C}^{y}_{jl})=0$). The GWOT component is able to tackle order variations and repeated actions [71].
Efficient and Robust Solution. Following [53], we can compute the GWOT component efficiently as $\mathcal{F}_{\text{GWOT}}(\mathbf{C}^{x},\mathbf{C}^{y},\mathbf{T})=\langle\mathbf{C}^{x}\mathbf{T}\mathbf{C}^{y},\mathbf{T}\rangle$. By adding an entropy regularization term $-\epsilon H(\mathbf{T})$, with $H(\mathbf{T})=-\sum_{i,j}T_{ij}\log T_{ij}$ and $\epsilon>0$, to the FGWOT formulation in Eq. 3, the solution $\mathbf{T}^{\star}$ can be derived efficiently, following [53]. Our solver often needs less than $25$ iterations to converge, and thanks to the sparse $\mathbf{C}^{x}$ and $\mathbf{C}^{y}$, each iteration has $O(NM)$ time complexity. In addition, to enhance robustness, we follow [47] to append a virtual frame to $X$ and $Y$ for tackling background/redundant frames. In particular, we include an additional row and column to $\mathbf{T}$ and expand other variables accordingly. If the matching probability of $x_{i}$ ($i\leq N$) with every $y_{j}$ ($j\leq M$) is below a threshold parameter $\zeta$, we assign $x_{i}$ to the virtual frame $y_{M+1}$, and vice versa. Note that virtual frames and their corresponding frames are not included in computing the losses.

### v2

@@ Abstract
We study self-supervised procedure learning, which discovers key steps and their order from a set of unlabeled videos. Previous methods typically learn frame-to-frame correspondences between videos before determining key steps and their order. However, their performance often suffers from order variations, background/redundant frames, and repeated actions. To overcome these challenges, we propose a self-supervised framework, which utilizes a fused Gromov-Wasserstein optimal transport with a structural prior for frame-to-frame mapping. However, optimizing only for the above temporal alignment may lead to degenerate solutions, where all frames are mapped to a small cluster in the embedding space and thus every video is assigned to just one key step. To address that issue, we integrate a contrastive regularization, which maps different frames to various points, avoiding trivial solutions. Finally, extensive experiments on egocentric and third-person benchmarks demonstrate our superior performance over prior works, including OPEL which relies on a classical Kantorovich optimal transport with an optimality prior.
@@ 1 Introduction
Understanding how to perform complex tasks by watching others is a hallmark of human intelligence. An increasingly important goal in computer vision is to replicate this ability in machines, enabling them to learn a procedure from watching video demonstrations. Procedure learning (PL) refers to the task of identifying key steps and their ordering from instructional videos [alayrac2016unsupervised, kukleva2019unsupervised, zhukov2019cross]. Unlike action recognition and segmentation, which focus on analyzing single videos and classifying short-term activities [simonyan2014two, carreira2017quo, piergiovanni2017learning], procedure learning extracts semantics such as key steps and their order across videos, even when step appearances vary, their order differ, or background/redundant frames occur frequently, as shown in Fig. 1. These challenges are especially common in real-world datasets.
@@
While earlier methods relied on full or weak supervision [bojanowski2015weakly, huang2016connectionist] which require substantial human effort and do not scale to large task domains, recent works have focused on self-supervised alignment of instructional videos. CnC [bansal2022my], for instance, uses temporal cycle-consistency and contrastive learning to align steps across videos. However, it is sensitive to step order variation and background/redundant frames. More recent works have employed optimal transport (OT) formulations for alignment due to their flexibility. One such method, VAVA [liu2022learning], considers alignment as an OT problem to accommodate variations in step ordering and incorporates an optimality prior along with inter-video and intra-video contrastive losses to solve the temporal misalignment problem. However, a key limitation in VAVA [liu2022learning] is that it is challenging to balance multiple losses and handle repeated actions, as noted by [donahue2024learning, ali2025joint]. Recently, OPEL [chowdhuryopel] applies VAVA for self-supervised procedure learning, inheriting these limitations.
@@
In this paper, we propose a self-supervised procedure learning framework built upon a fused Gromov-Wasserstein optimal transport (FGWOT) formulation with a structural prior. Unlike previous OT-based approaches [liu2022learning, chowdhuryopel] that utilize a traditional Kantorovich OT formulation, our method not only aligns frames based on appearance but also enforces a structural prior (i.e., temporal consistency) across videos. This makes our model better suited for handling step order variations, background/redundant frames, and repeated actions. Moreover, we empirically identify that optimizing only for temporal alignment may lead to degenerate solutions, where all frames collapse to a single cluster in the embedding space. To prevent this issue, we utilize Contrastive Inverse Difference Moment (C-IDM) as a regularization that encourages embedding diversity across frames. Our regularized Gromov-Wasserstein optimal transport (RGWOT) approach uses a unified loss with a purpose-aligned regularization, avoiding the difficulty of balancing multiple losses and conflicting regularizations. Our model outperforms prior works on both egocentric [bansal2022my] and third-person [elhamifar2019unsupervised, zhukov2019cross] procedure learning benchmarks..
@@
  • We introduce an optimal transport-based framework for self-supervised procedure learning, which adopts a fused Gromov-Wasserstein optimal transport formulation with a structural prior for frame-to-frame correspondences.
@@
  • Extensive evaluations on egocentric and third-person datasets show that our approach outperforms prior works, including OPEL which adopts a classical Kantorovich optimal transport formulation with an optimality prior.
@@ 2 Related Work
Self-Supervised Learning. Self-supervised learning uses pretext tasks that generate a supervision signal from the data itself. Early self-supervised approaches focused mainly on images, leveraging spatial cues through tasks such as image colorization [larsson2016learning, larsson2017colorization], object counting [liu2018leveraging, noroozi2017representation], solving puzzles [carlucci2019domain, kim2019self, kim2018learning], predicting rotations [feng2019self, gidaris2018unsupervised], and image inpainting [jenni2020steering]. Recently, self-supervised learning has expanded to video data, which inherently offers spatial and temporal structures. Video-based pretext tasks include forecasting future frames [ahsan2018discrimnet, diba2019dynamonet, srivastava2015unsupervised, vondrick2016generating, han2019video], enforcing temporal coherence [goroshin2015unsupervised, mobahi2009deep, zou2011unsupervised], predicting the temporal order of frames [fernando2017self, lee2017unsupervised, misra2016shuffle, xu2019self], determining the arrow of time [pickup2014seeing, wei2018learning], and predicting the pace of actions [benaim2020speednet, wang2020self, yao2020video], and frame clustering [kumar2022unsupervised, tran2024permutation]. In this work, we focus on self-supervised procedure learning for identifying key steps and their order from videos.
@@
Procedure Learning. Many procedure learning (PL) methods focus on learning frame-level features that capture task structure [elhamifar2019unsupervised, kukleva2019unsupervised, elhamifar2020self, vidalmata2021joint, bansal2022my]. For example, Kukleva et al. [kukleva2019unsupervised] learn frame representation by using timestamps, while VidalMata et al. [vidalmata2021joint] focus on predicting future frames. Elhamifar et al. [elhamifar2020self] utilize attention mechanism on individual frames for better feature learning. Bansal et al. [bansal2022my] use temporal correspondences across videos to create signals for robust frame-level embeddings. Recently, Bansal et al. [bansal2024united] use a task-level graph representation to cluster frames that are semantically similar and temporally close. Beyond purely visual approaches, procedure learning has also been explored in multi-modal settings, combining narrated text and videos [damen2014you, yu2014instructional, malmaud2015s, alayrac2016unsupervised, doughty2020action, fried2020learning], as well as modalities like optical flow, depth, and gaze information [shah2023steps]. These multi-modal approaches generally assume reliable alignment between modalities [alayrac2016unsupervised, malmaud2015s, yu2014instructional], an assumption that frequently fails due to asynchrony among data streams [elhamifar2019unsupervised, elhamifar2020self]. Also, reliance on imperfect automatic speech recognition requires manual correction, and multi-modal integration adds memory/computation. A recent PL work [chowdhuryopel] focuses purely on visual data, bypassing the inaccuracies associated with multi-modal alignment. Here, we propose an advanced optimal transport formulation, outperforming [chowdhuryopel] on visual procedure learning.
@@
Video Alignment. TCC [dwibedi2019tcc] enforces cycle consistency, while GTCC [donahue2024learning] extends it with multi-cycle consistency to handle repeated actions. However, they focus on local alignment and do not capture the global temporal structure. For global alignment, LAV [haresh2021learning] applies dynamic time warping [cuturi2017soft] but assumes monotonic ordering without background/redundant frames. VAVA [liu2022learning] addresses non-monotonic sequences and background/redundant frames using a classical Kantorovich optimal transport [cuturi2013sinkhorn] but suffers from balancing multiple losses and repeated actions. Recently, VAOT [ali2025joint] develops a fused Gromov-Wasserstein optimal transport which can handle order variations, background/redundant frames, and repeated actions. We empirically show a possible degenerate solution for VAOT [ali2025joint] and propose a regularized optimal transport to address that.
@@
Learning Key-Step Ordering. People often perform the same task using different orders of key-steps, but most prior works fail to model this variation. They either assume a strict step ordering [elhamifar2019unsupervised, kukleva2019unsupervised, vidalmata2021joint] or do not predict the order at all [elhamifar2020self, shen2021learning]. To better capture how tasks are actually completed, it is important to account for these different key-step sequences. Our method addresses this by creating a specific sequence of key-steps for each video, allowing the model to infer the particular order required to complete the task.
Temporal Contrastive Learning. Inspired by the success of contrastive learning for self-supervised image representation learning [chen2020simple, he2020momentum, caron2020unsupervised], temporal contrastive learning has been proposed for video representation learning [qian2021spatiotemporal, basu2022unsupervised, dave2022tclr]. For example, TCLR [dave2022tclr] applies inter- and intra-instance temporal contrastive losses to capture temporal variations between and within videos. Unlike these methods which use contrastive learning as their main pretext task, we use video alignment as our main pretext task, and contrastive learning as our regularization to avoid trivial solutions.
@@ 3 Our Approach
Our approach includes finding frame-to-frame correspondences between videos in Sec. 3.1 and performing clustering and key-step ordering in Sec. 3.2.
Notations. Let us represent two input videos of $N$ and $M$ frames as $X=\{x_{1},\dots,x_{N}\}$ and $Y=\{y_{1},\dots,y_{M}\}$ respectively. The frame-level embeddings of $X$ and $Y$ are then written as $\mathbf{X}=f_{\bm{\theta}}(X)\in\mathbb{R}^{N\times D}$ and $\mathbf{Y}=f_{\bm{\theta}}(Y)\in\mathbb{R}^{M\times D}$ respectively, where $f_{\bm{\theta}}$ (with learnable parameters $\bm{\theta}$) is the embedding function and $D$ is the length of the embedding vector. $K$ is the number of key steps. In addition, the dot product of $\mathbf{A},\mathbf{B}\in\mathbb{R}^{n\times m}$ is expressed as $\langle\mathbf{A},\mathbf{B}\rangle=\sum_{i,j}A_{ij}B_{ij}$, while a vector of ones is written as $\mathbf{1}_{n}\in\mathbb{R}^{n}$. Finally, $n]=\{1,\dots,n\}$ models a discrete set of $n$ elements. We represent the (m1)(m-1) dimensional probability simplex as $\Delta_{m}\subset\mathbb{R}^{m}$ and the Cartesian product space consisting of $n$ such simplexes as $\Delta_{m}^{n}\subset\mathbb{R}^{m\times n}$.
@@ 3.1.1 Fused Gromov-Wasserstein Optimal Transport
We consider temporal video alignment as an optimal transport problem in the discrete setting. We follow [ali2025joint] to apply a fused Gromov-Wasserstein optimal transport formulation (FGWOT), which is a combination of Kantorovich optimal transport (KOT) and Gromov-Wasserstein optimal transport (GWOT). Below we provide the details of KOT and GWOT. In particular, given a ground cost $\mathbf{C}\in\mathbb{R}^{n\times m}_{+}$, the KOT objective seeks the minimum-cost coupling $\mathbf{T}^{\star}$ between histograms $\mathbf{p}\in\Delta_{n}$ and $\mathbf{q}\in\Delta_{m}$ as:
@@
Cost Matrices. The cost matrix $\mathbf{C}\in\mathbb{R}^{N\times M}_{+}$ in the KOT component measures visual differences between $X$ and $Y$ and is defined as $\mathbf{C}_{ij}=1-\frac{\mathbf{x}_{i}^{\top}\mathbf{y}_{j}}{\|\mathbf{x}_{i}\|_{2}\|\mathbf{y}_{j}\|_{2}}$, with frame embeddings $\mathbf{x}_{i}=f_{\bm{\theta}}(x_{i})$ and $\mathbf{y}_{j}=f_{\bm{\theta}}(y_{j})$. Also, we follow [liu2022learning] to augment $\mathbf{C}$ as $\tilde{\mathbf{C}}=\mathbf{C}+\rho\mathbf{R}$, where $\rho\geq 0$ is a balancing parameter and the temporal prior $\mathbf{R}$ (with $\mathbf{R}_{ij}=\left|i/N-j/M\right|$) enforces the coupling $\mathbf{T}$ to have a banded diagonal shape. Moreover, we set $L(a,b)=ab$, and follow [ali2025joint] to define cost matrices $\mathbf{C}^{x}\in\mathbb{R}_{+}^{N\times N}$ and $\mathbf{C}^{y}\in\mathbb{R}_{+}^{M\times M}$ in the GWOT component as:
@@
with $\delta_{ik}=|i-k|$, $\delta_{jl}=|j-l|$, and a radius parameter $r\in(0,1]$. The GWOT component imposes temporal consistency over $\mathbf{T}$. Specifically, mapping temporally adjacent frames in $X$ ($\delta_{ik}\leq Nr$) to temporally remote frames in $Y$ ($\delta_{jl}>Mr$) induces a cost ($L(\mathbf{C}^{x}_{ik},\mathbf{C}^{y}_{jl})=\frac{1}{r}$), however assigning temporally neighboring frames in $X$ ($\delta_{ik}\leq Nr$) to temporally nearby frames in $Y$ ($\delta_{jl}\leq Mr$) or assigning temporally remote frames in $X$ ($\delta_{ik}>Nr$) to temporally distant frames in $Y$ ($\delta_{jl}>Mr$) induces no cost ($L(\mathbf{C}^{x}_{ik},\mathbf{C}^{y}_{jl})=0$). The GWOT component is able to tackle order variations and repeated actions [xu2024temporally].
Efficient and Robust Solution. Following [peyre2016gromov], we can compute the GWOT component efficiently as $\mathcal{F}_{\text{GWOT}}(\mathbf{C}^{x},\mathbf{C}^{y},\mathbf{T})=\langle\mathbf{C}^{x}\mathbf{T}\mathbf{C}^{y},\mathbf{T}\rangle$. By adding an entropy regularization term $\epsilon H(\mathbf{T})$, with $H(\mathbf{T})=-\sum_{i,j}T_{ij}\log T_{ij}$ and $\epsilon>0$, to the FGWOT formulation in Eq. 3, the solution $\mathbf{T}^{\star}$ can be derived efficiently, following [peyre2016gromov]. Our solver often needs less than $25$ iterations to converge, and thanks to the sparse $\mathbf{C}^{x}$ and $\mathbf{C}^{y}$, each iteration has $O(NM)$ time complexity. In addition, to enhance robustness, we follow [liu2022learning] to append a virtual frame to $X$ and $Y$ for tackling background/redundant frames. In particular, we include an additional row and column to $\mathbf{T}$ and expand other variables accordingly. If the matching probability of $x_{i}$ ($i\leq N$) with every $y_{j}$ ($j\leq M$) is below a threshold parameter $\zeta$, we assign $x_{i}$ to the virtual frame $y_{M+1}$, and vice versa. Note that virtual frames and their corresponding frames are not included in computing the losses.